# Interpolation in Classical Propositional Logic

### v1

@@ 7 Interpolants via Quantifier Elimination
The proof has two steps: (i) we show that uniform interpolants always exist in the language obtained from propositional logic by adding quantifiers over atoms; and (ii) we show that these quantifiers can be always eliminated in sense that one can always construct a logically equivalent propositional formula that does not use quantifiers. Overall, one then obtains that uniform interpolants always exist.
@@
Validity, satisfiability, and logical equivalence are now defined in the obvious way. The definition of the signature $\text{sig}(\varphi)$ of a QBF $\varphi$ extends inductively the definition for propositional logic with $\text{sig}(\exists p.\varphi)=\text{sig}(\varphi)\setminus\{p\}$.
@@
QBF Uniform Interpolants. Assume $\varphi$ and a signature $\sigma\subseteq\text{sig}(\varphi)$ are given. Let $p_{1},\ldots,p_{n}$ be any ordering of the atoms in $\mathit{sig}(\varphi)\setminus\sigma$ and consider the QBF $\chi=\exists p_{1}\cdots\exists p_{n}.\varphi$. Then
  • $\varphi\models\chi$ and $\text{sig}(\chi)=\sigma$;
  • for every propositional formula $\psi$, if $\varphi\models\psi$ and $\text{sig}(\psi)\cap\mathit{sig}(\varphi)\subseteq\sigma$, then $\chi\models\psi$.
@@
We conclude this section with a few further observations. Let $\forall p.\varphi:=\neg\exists p.\neg\varphi$. When constructing Craig interpolants for formulas $\varphi,\psi$, instead of $\exists p_{1}\cdots\exists p_{n}.\varphi$ with $p_{1},\ldots,p_{n}$ any ordering of $\text{sig}(\varphi)\setminus\text{sig}(\psi)$, we could have also used $\forall q_{1}\cdots\forall q_{n}.\psi$ with $q_{1},\ldots,q_{n}$ any ordering of $\text{sig}(\psi)\setminus\text{sig}(\varphi)$. The formula $\exists p_{1}\cdots\exists p_{n}.\varphi$ is (up to equivalence) the logically strongest interpolant for $\varphi,\psi$ (it entails all others) and the formula $\forall q_{1}\cdots\forall q_{n}.\psi$ is (up to equivalence) the logically weakest interpolant for $\varphi,\psi$ (it is entailed by all others).
@@ 8 Interpolants via DNF
The uniform $\sigma$-interpolant for a formula $\varphi$ formalises a natural notion of forgetting the atoms in $\mathit{sig}(\varphi)\setminus\sigma$ from $\varphi$. This suggests that one can construct the uniform $\sigma$-interpolant by somehow dropping the atoms in $\mathit{sig}(\varphi)\setminus\sigma$ from $\varphi$ in a syntactic sense. This is clearly not the case in general. For instance, for $\varphi=(p\vee q)\wedge\neg p$ and $\sigma=\{p\}$, we do not obtain the uniform $\sigma$-interpolant for $\varphi$ by just dropping $q$: the result of dropping $q$ is logically equivalent to $p\wedge\neg p$ and $\varphi\not\models p\wedge\neg p$. If, however, $\varphi$ is in DNF, then we indeed obtain the uniform $\sigma$-interpolant by dropping the atoms in $\sigma$.
@@
with $\ell_{ij}$ literals. Let $\sigma\subseteq\text{sig}(\varphi)$. We may assume that no conjunctive clause $\varphi_{i}$ contains an atom and its negation. Then obtain $\varphi^{-\sigma}$ from $\varphi$ by dropping all literals $\ell_{ij}$ not in $\sigma$ from every $\varphi_{i}$.
@@
$\varphi=\bigvee_{1\leq i\leq n}\varphi_{i},\quad\varphi_{i}=\bigwedge_{1\leq j\leq n_{i}}\ell_{ij},\quad\psi=\bigvee_{1\leq i^{\prime}\leq m}\psi_{i^{\prime}},\quad\psi_{i^{\prime}}=\bigwedge_{1\leq j^{\prime}\leq m_{i}}\ell_{i^{\prime}j^{\prime}}^{\prime}$
and that $\varphi\wedge\psi$ is not satisfiable. We may assume that no $\varphi_{i}$ nor $\psi_{i}$ contains an atom and its negation. We construct a Craig separator as follows. Let $\sigma=\text{sig}(\varphi)\cap\text{sig}(\psi)$. Because $\varphi\wedge\psi$ is not satisfiable, for any pair $i,i^{\prime}$ we find $j,j^{\prime}$ s.t. $\ell_{ij}\wedge\ell_{i^{\prime}j^{\prime}}^{\prime}$ unsatisfiable. Take any collection $P$ of such pairs $(\ell_{ij},\ell_{i^{\prime}j^{\prime}}^{\prime})$. Then $P$ can be regarded as a proof that $\varphi\wedge\psi$ is not satisfiable. Observe that each pair in $P$ consists of an atom in $\sigma$ and its negation. The formula $\varphi_{P}$ now is the disjunction over all literals from $\varphi$ that contribute to a pair in $P$. It is easy to check that $\varphi_{P}$ satisfies all properties of a Craig separator: $\varphi\models\varphi_{P}$, $\text{sig}(\varphi_{P})\subseteq\sigma$, and $\varphi_{P}\wedge\psi$ is unsatisfiable since the pairs in $P$ still show that $\varphi_{P}\wedge\psi$ is unsatisfiable. Indeed, it is even a Craig-Lyndon separator: Any literal in $\varphi_{P}$ occurs as a literal $\ell_{ij}$ in $\varphi$, while its complement occurs as a literal $\ell_{i^{\prime}j^{\prime}}^{\prime}$ in $\psi$. We obtain the following theorem, which is at the same time our first proof of Theorem˜3.8, namely that propositional logic has the LIP.
@@ 9 Interpolants via Resolution
Resolution for propositional logic is a decision procedure for propositional logic that operates on formulas in CNF, which in this context are seen as sets of clauses. In the case of propositional logic, the central inference rule is the resolution rule:
@@ 9.1 Uniform Interpolation Using Resolution
Variant 1: Using quantifier elimination. Assume that no clause contains both $p$ and $\neg p$. (Since such clauses are tautologies, we can safely remove them.) By factoring out occurrences of $p$ and $\neg p$ respectively, we can transform $\Phi$ into an equivalent formula of the following form, where $\Phi_{1}$, $\Phi_{2}$ and $\Phi_{3}$ are formulas in CNF that contain no occurrences of $p$:
@@
Variant 2: Using ordered resolution. For this proof, we use refutational completeness of ordered resolution, where we use an ordering $\prec_{p}$ that makes $p$ maximal. We need to show that $\Psi$ is a uniform $\Sigma$-interpolant of $\Phi$ for $\Sigma=\mathit{sig}(\Phi)\setminus\{p\}$. The first two conditions in Definition˜4.10 hold because 1) the resolvent is always entailed by its parent clauses, and 2) we explicitly removed all occurrences of $p$. It remains to show that for any formula $\varphi$ that does not contain $p$, $\Psi\models\varphi$ if $\Phi\models\varphi$. Since $\Phi\models\varphi$ iff $\Phi\wedge\neg\varphi\models\bot$, we can do so by showing that for any formula $\Theta$ in CNF s.t. $p\not\in\mathit{sig}(\Theta)$, $\Phi\cup\Theta\models\bot$ implies $\Psi\cup\Theta\models\bot$. We can decide $\Phi\cup\Theta\models\bot$ using ordered resolution with the ordering $\prec_{p}$. In particular, this means that we first compute all resolvents upon $p$ (which only occurs in $\Phi$, but not in $\Theta$), and then ignore occurrences of $p$ while we continue. Effectively, we thus first compute $\Psi$, and then continue on $\Psi\cup\Theta$. Thus, if the empty clause can be derived from $\Phi\cup\Theta$, it can also be derived from $\Psi\cup\Theta$.
@@ 9.2 Craig Interpolation Using Resolution
By induction of the steps in such a proof, we obtain that each clause is annotated whose size is polynomial in the number of proof steps that lead to this clause. For the input clauses, this holds trivially. By analyzing the different cases for how we annotate new resolvents, we see that each intermediate separator introduces at most five new subformulas. Consequently, the (dag-)size of the computed separator is at most five times the number of inference steps, and thus linearly bounded in the size of the proof.
@@
10 Craig Interpolants via Tableaux
@@ 10.2 Craig-Lyndon Separator Extraction from a Closed Tableau
Lemma 10.42 is easy to verify. For example, Lemma 10.42 expresses that, for arbitrary sets $S$ of biased formulas, the literal $\ell$ is a Craig-Lyndon separator for $\bigwedge_{L(\theta)\in S}\theta\land\ell,\;\bigwedge_{R(\theta)\in S}\theta\land\overline{\ell}$. This holds, since $\bigwedge_{L(\theta)\in S}\theta\land\ell\models\ell$, $\bigwedge_{R(\theta)\in S}\theta\land\overline{\ell}\land\ell\models\bot$, and $\mathit{sig}^{\pm}(\ell)=\{\ell\}\subseteq\mathit{sig}^{\pm}(\bigwedge_{L(\theta)\in S}\theta\land\ell)\cap\mathit{sig}^{\pm}(\lnot(\bigwedge_{R(\theta)\in S}\theta\land\overline{\ell}))$.
@@ 11 The Size of Interpolants
with $q$ ranging over atoms in $\varphi_{n}$, and $\psi\wedge\chi$, $\psi\vee\chi$, and $\neg\psi$ ranging over subformulas of $\varphi_{n}$. Let $\sigma_{n}=\text{sig}(\varphi_{n})$. Then $\psi_{n}\models p_{\varphi_{n}}\leftrightarrow\varphi_{n}$ and so we have explicit $\sigma_{n}$-definitions of $p_{\varphi_{n}}$ under $\psi_{n}$. There cannot be polynomial tree-size explicit $\sigma_{n}$-definitions of $p_{\varphi_{n}}$ under $\psi_{n}$ since these would be propositional formulas equivalent to $\varphi_{n}$ of polynomial tree-size. It follows from the proof of Theorem 5.16 that Craig interpolants can also not be of polynomial tree-size.
In what follows, we only consider the dag-size of a formula. So we say that propositional logic has polysize Craig interpolants if there is a polynomial function $f:\mathbb{N}\times\mathbb{N}\rightarrow\mathbb{N}$ such that for any $\varphi,\psi$ with $\varphi\models\psi$ there exists a Craig interpolant $\chi$ for $\varphi,\psi$ with $|\chi|\leq f(|\varphi|,|\psi|)$. We say that propositional logic has polysize uniform interpolants if there is a polynomial function $f:\mathbb{N}\rightarrow\mathbb{N}$ such that for any $\varphi$ and ordering ${\bf p}=p_{1},\ldots,p_{n}$ of any subset of $\text{sig}(\varphi)$, there exists a propositional formula $\chi$ equivalent to $\exists{\bf p}.\varphi$ with $|\chi|\leq f(|\varphi|)$, where $\exists{\bf p}.\varphi$ stands for $\exists p_{1}\cdots\exists p_{n}.\varphi$.
@@
Conversely, we show that ${\bf NP}\subseteq{\bf P}_{{\bf/poly}}$ implies that uniform interpolants of polynomial size exist. Let $\exists P_{n}$ denote the set of formulas of the form $\varphi=\exists q_{1}\cdots\exists q_{m}.\psi$ with $\psi$ a propositional formula and $\text{sig}(\varphi)=\{p_{1},\ldots,p_{n}\}$ the remaining atoms in $\varphi$ distinct from $q_{1},\ldots,q_{m}$. It suffices to show that ${\bf NP}\subseteq{\bf P}_{{\bf/poly}}$ implies that formulas in $\exists P=\bigcup_{n\in\mathbb{N}}\exists P_{n}$ are equivalent to polysize propositional formulas. Let $w_{\varphi}\in\{0,1\}^{\ast}$ be an encoding of formulas $\varphi\in\exists P$. We regard every $w\in\{0,1\}^{n}$ as a model with $w(p_{i})=w_{i}$ for $1\leq i\leq n$. Consider the set
@@
12 Bibliographic and Historic Remarks
12.1 Interpolation via Resolution
Resolution was introduced in 1965 by J. Alan Robinson [41] as a machine-suited proof method for first-order logic. For propositional logic, it was, in the dual form for DNF, already established since 1955 as the consensus method by Willard Van Orman Quine [40], which is applied to convert a given DNF to a form that consists of all its prime implicants (minimal conjunctive clauses without complementary literals that imply the overall formula). The first appearance of propositional resolvent/consensus construction is attributed to Archie Blake in 1937 [6, 9]. Our resolution-based method for uniform interpolation from Section 9.1 has descendants a in more general method for second-order quantifier elimination on the basis of first-order logic developed by Wilhelm Ackermann in 1935 [1], which was rediscovered in 1992 by Dov M. Gabbay and Hans Jürgen Ohlbach [16]. In its dual variation, our method for propositional logic is attributed to Edward W. Samson and Rolf K. Mueller in 1955 [45] (see [27, Sect. 7.2.2.2]). They address the tautology problem by eliminating all atoms from a given propositional DNF one by one with consensus. The system for Craig interpolation shown in Section 9.2 is called HKPYM by Maria Paola Bonacina and Moa Johansson in their survey on ground interpolation systems [8], reflecting the initials of several authors who apparently discovered it independently, Guoxiang Huang in 1995 [21] (actually for first-order clauses), Jan Krajícek [31] in 1997, and Pavel Pudlák [39] also in 1997, as well as Greta Yorsh and Madanlal Musuvathi [52] for contributing in 2005 a completeness proof. In their survey, Bonacina and Johansson also present a second resolution-based propositional interpolation system by Kenneth L. McMillan from 2003 [33, 35] in a form that allows easy comparison.
12.2 Interpolation via Tableaux
Analytic tableaux became widely known with Raymond Smullyan’s compact monograph on first-order logic from 1968 [47]. They are related to earlier systems by, among others, Jaakko Hintikka 1955 [20] and Beth 1959 [5]. For a comprehensive historic account, see Melvin Fitting’s introduction [15] to the Handbook of Tableau Methods [13]. A monograph by Fitting [14] and a handbook article by Reiner Hähnle [22] give presentations of tableaux that take aspects of automated reasoning into account. The interpolant extraction from analytic tableaux in Section 10 is based on the monograph by Fitting from 1996 [14] and the earlier monograph by Smullyan [47]. Fitting specifies it by means of a rule system. Similar interpolation methods were presented on the basis of semantic tableaux by Beth in 1959 [5] and on the basis of sequent calculi by Shoji Maehara in 1960 [32, 49] and also by Smullyan in 1968 [47].

### v2

@@ 7 Interpolants via Quantifier Elimination
The proof has two steps: (i) we show that uniform interpolants always exist in the language obtained from propositional logic by adding quantifiers over atoms; and (ii) we show that these quantifiers can be always eliminated in the sense that one can always construct a logically equivalent propositional formula that does not use quantifiers. Overall, one then obtains that uniform interpolants always exist.
@@
Validity, satisfiability, and logical equivalence are now defined in the obvious way. The definition of the signature $\mathit{sig}(\varphi)$ of a QBF $\varphi$ extends inductively the definition for propositional logic with $\mathit{sig}(\exists p.\varphi)=\mathit{sig}(\varphi)\setminus\{p\}$.
@@
QBF Uniform Interpolants. Assume $\varphi$ and a signature $\sigma\subseteq\mathit{sig}(\varphi)$ are given. Let $p_{1},\ldots,p_{n}$ be any ordering of the atoms in $\mathit{sig}(\varphi)\setminus\sigma$ and consider the QBF $\chi=\exists p_{1}\cdots\exists p_{n}.\varphi$. Then
  • $\varphi\models\chi$ and $\mathit{sig}(\chi)=\sigma$;
  • for every propositional formula $\psi$, if $\varphi\models\psi$ and $\mathit{sig}(\psi)\cap\mathit{sig}(\varphi)\subseteq\sigma$, then $\chi\models\psi$.
@@
We conclude this section with a few further observations. Let $\forall p.\varphi:=\neg\exists p.\neg\varphi$. When constructing Craig interpolants for formulas $\varphi,\psi$, instead of $\exists p_{1}\cdots\exists p_{n}.\varphi$ with $p_{1},\ldots,p_{n}$ any ordering of $\mathit{sig}(\varphi)\setminus\mathit{sig}(\psi)$, we could have also used $\forall q_{1}\cdots\forall q_{n}.\psi$ with $q_{1},\ldots,q_{n}$ any ordering of $\mathit{sig}(\psi)\setminus\mathit{sig}(\varphi)$. The formula $\exists p_{1}\cdots\exists p_{n}.\varphi$ is (up to equivalence) the logically strongest interpolant for $\varphi,\psi$ (it entails all others) and the formula $\forall q_{1}\cdots\forall q_{n}.\psi$ is (up to equivalence) the logically weakest interpolant for $\varphi,\psi$ (it is entailed by all others).
@@ 8 Interpolants via DNF
The uniform $\sigma$-interpolant for a formula $\varphi$ formalises a natural notion of forgetting the atoms in $\mathit{sig}(\varphi)\setminus\sigma$ from $\varphi$. This suggests that one can construct the uniform $\sigma$-interpolant by somehow dropping the atoms in $\mathit{sig}(\varphi)\setminus\sigma$ from $\varphi$ in a syntactic sense. This is clearly not the case in general. For instance, for $\varphi=(p\vee q)\wedge\neg p$ and $\sigma=\{p\}$, we do not obtain the uniform $\sigma$-interpolant for $\varphi$ by just dropping $q$: the result of dropping $q$ is logically equivalent to $p\wedge\neg p$ and $\varphi\not\models p\wedge\neg p$. If, however, $\varphi$ is in DNF, then we indeed obtain the uniform $\sigma$-interpolant by dropping the atoms in $\mathit{sig}(\varphi)\setminus\sigma$.
@@
with $\ell_{ij}$ literals. Let $\sigma\subseteq\mathit{sig}(\varphi)$. We may assume that no conjunctive clause $\varphi_{i}$ contains an atom and its negation. Then obtain $\varphi^{-\sigma}$ from $\varphi$ by dropping all literals $\ell_{ij}$ not in $\sigma$ from every $\varphi_{i}$.
@@
$\varphi=\bigvee_{1\leq i\leq n}\varphi_{i},\quad\varphi_{i}=\bigwedge_{1\leq j\leq n_{i}}\ell_{ij},\quad\psi=\bigvee_{1\leq i^{\prime}\leq m}\psi_{i^{\prime}},\quad\psi_{i^{\prime}}=\bigwedge_{1\leq j^{\prime}\leq m_{i^{\prime}}}\ell_{i^{\prime}j^{\prime}}^{\prime}$
and that $\varphi\wedge\psi$ is not satisfiable. We may assume that no $\varphi_{i}$ nor $\psi_{i}$ contains an atom and its negation. We construct a Craig separator as follows. Let $\sigma=\mathit{sig}(\varphi)\cap\mathit{sig}(\psi)$. Because $\varphi\wedge\psi$ is not satisfiable, for any pair $i,i^{\prime}$ we find $j,j^{\prime}$ s.t. $\ell_{ij}\wedge\ell_{i^{\prime}j^{\prime}}^{\prime}$ unsatisfiable. Take any collection $P$ of such pairs $(\ell_{ij},\ell_{i^{\prime}j^{\prime}}^{\prime})$. Then $P$ can be regarded as a proof that $\varphi\wedge\psi$ is not satisfiable. Observe that each pair in $P$ consists of an atom in $\sigma$ and its negation. The formula $\varphi_{P}$ now is the disjunction over all literals from $\varphi$ that contribute to a pair in $P$. It is easy to check that $\varphi_{P}$ satisfies all properties of a Craig separator: $\varphi\models\varphi_{P}$, $\mathit{sig}(\varphi_{P})\subseteq\sigma$, and $\varphi_{P}\wedge\psi$ is unsatisfiable since the pairs in $P$ still show that $\varphi_{P}\wedge\psi$ is unsatisfiable. Indeed, it is even a Craig-Lyndon separator: Any literal in $\varphi_{P}$ occurs as a literal $\ell_{ij}$ in $\varphi$, while its complement occurs as a literal $\ell_{i^{\prime}j^{\prime}}^{\prime}$ in $\psi$. We obtain the following theorem, which is at the same time our first proof of Theorem˜3.8, namely that propositional logic has the LIP.
@@ 9 Interpolants via Resolution
Resolution for propositional logic operates on formulas in CNF, which in this context are seen as sets of clauses. Its central inference rule is the resolution rule:
@@ 9.1 Uniform Interpolation Using Resolution
Variant 1: Using quantifier elimination.
Assume that no clause contains both $p$ and $\neg p$. (Since such clauses are tautologies, we can safely remove them.) By factoring out occurrences of $p$ and $\neg p$ respectively, we can transform $\Phi$ into an equivalent formula of the following form, where $\Phi_{1}$, $\Phi_{2}$ and $\Phi_{3}$ are formulas in CNF that contain no occurrences of $p$:
@@
Variant 2: Using ordered resolution.
For this proof, we use refutational completeness of ordered resolution, where we use an ordering $\prec_{p}$ that makes $p$ maximal. We need to show that $\Psi$ is a uniform $\Sigma$-interpolant of $\Phi$ for $\Sigma=\mathit{sig}(\Phi)\setminus\{p\}$. The first two conditions in Definition˜4.10 hold because 1) the resolvent is always entailed by its parent clauses, and 2) we explicitly removed all occurrences of $p$. It remains to show that for any formula $\varphi$ that does not contain $p$, $\Psi\models\varphi$ if $\Phi\models\varphi$. Since $\Phi\models\varphi$ iff $\Phi\wedge\neg\varphi\models\bot$, we can do so by showing that for any formula $\Theta$ in CNF s.t. $p\not\in\mathit{sig}(\Theta)$, $\Phi\cup\Theta\models\bot$ implies $\Psi\cup\Theta\models\bot$. We can decide $\Phi\cup\Theta\models\bot$ using ordered resolution with the ordering $\prec_{p}$. In particular, this means that we first compute all resolvents upon $p$ (which only occurs in $\Phi$, but not in $\Theta$), and then ignore occurrences of $p$ while we continue. Effectively, we thus first compute $\Psi$, and then continue on $\Psi\cup\Theta$. Thus, if the empty clause can be derived from $\Phi\cup\Theta$, it can also be derived from $\Psi\cup\Theta$.
@@ 9.2 Craig Interpolation Using Resolution
By induction on the number of steps in the proof, we obtain that each clause is annotated by a formula whose size is polynomial in the number of proof steps that lead to this clause. For the input clauses, this holds trivially. By analyzing the different cases for how we annotate new resolvents, we see that each intermediate separator introduces at most five new subformulas. Consequently, the (dag-)size of the computed separator is at most five times the number of inference steps, and thus linearly bounded in the size of the proof.
@@
9.3 Bibliographic and Historic Remarks
Resolution was introduced in 1965 by J. Alan Robinson [41] as a machine-suited proof method for first-order logic. For propositional logic, it was, in the dual form for DNF, already established since 1955 as the consensus method by Willard Van Orman Quine [40], which is applied to convert a given DNF to a form that consists of all its prime implicants (minimal conjunctive clauses without complementary literals that imply the overall formula). The first appearance of propositional resolvent/consensus construction is attributed to Archie Blake in 1937 [6, 9]. Our resolution-based method for uniform interpolation from Section 9.1 has descendants in a more general method for second-order quantifier elimination on the basis of first-order logic developed by Wilhelm Ackermann in 1935 [1], which was rediscovered in 1992 by Dov M. Gabbay and Hans Jürgen Ohlbach [16]. In its dual variation, our method for propositional logic is attributed to Edward W. Samson and Rolf K. Mueller in 1955 [45] (see [27, Sect. 7.2.2.2]). They address the tautology problem by eliminating all atoms from a given propositional DNF one by one with consensus. The system for Craig interpolation shown in Section 9.2 is called HKPYM by Maria Paola Bonacina and Moa Johansson in their survey on ground interpolation systems [8], reflecting the initials of several authors who apparently discovered it independently, Guoxiang Huang in 1995 [22] (actually for first-order clauses), Jan Krajícek [31] in 1997, and Pavel Pudlák [39] also in 1997, as well as Greta Yorsh and Madanlal Musuvathi [52] for contributing in 2005 a completeness proof. In their survey, Bonacina and Johansson also present a second resolution-based propositional interpolation system by Kenneth L. McMillan from 2003 [33, 35] in a form that allows easy comparison.
10 Craig Interpolants via Analytic Tableaux
@@ 10.2 Craig-Lyndon Separator Extraction from a Closed Tableau
Lemma 10.42 is easy to verify. For example, Lemma 10.42 expresses that, for arbitrary sets $S$ of biased formulas, the literal $\ell$ is a Craig-Lyndon separator for $\bigwedge_{L(\theta)\in S}\theta\land\ell,\;\bigwedge_{R(\theta)\in S}\theta\land\overline{\ell}$. In case $\ell$ is a positive literal $p$, this holds, since $\bigwedge_{L(\theta)\in S}\theta\land p\models p$, $\bigwedge_{R(\theta)\in S}\theta\land\lnot p\land p\models\bot$, and $\mathit{sig}^{\pm}(p)=\{\langle p,{+}\rangle\}\subseteq\mathit{sig}^{\pm}(\bigwedge_{L(\theta)\in S}\theta\land p)\cap\mathit{sig}^{\pm}(\lnot(\bigwedge_{R(\theta)\in S}\theta\land\lnot p))$. The case for $\ell$ a negative literal $\lnot p$ holds analogously.
@@
10.3 Bibliographic and Historic Remarks
Analytic tableaux became widely known with Raymond Smullyan’s compact monograph on first-order logic from 1968 [47]. They are related to earlier systems by, among others, Jaakko Hintikka 1955 [21] and Evert W. Beth 1959 [5]. For a comprehensive historic account, see Melvin Fitting’s introduction [15] to the Handbook of Tableau Methods [13]. A monograph by Fitting [14] and a handbook article by Reiner Hähnle [18] give presentations of tableaux that take aspects of automated reasoning into account. Our interpolant extraction from analytic tableaux is based on the monograph by Fitting from 1996 [14] and the earlier monograph by Smullyan [47]. Fitting specifies it by means of a rule system. Similar interpolation methods were presented on the basis of semantic tableaux by Beth in 1959 [5] and on the basis of sequent calculi by Shoji Maehara in 1960 [32, 49] and also by Smullyan in 1968 [47].
@@ 11 The Size of Interpolants
with $q$ ranging over atoms in $\varphi_{n}$, and $\psi\wedge\chi$, $\psi\vee\chi$, and $\neg\psi$ ranging over subformulas of $\varphi_{n}$. Let $\sigma_{n}=\mathit{sig}(\varphi_{n})$. Then $\psi_{n}\models p_{\varphi_{n}}\leftrightarrow\varphi_{n}$ and so we have explicit $\sigma_{n}$-definitions of $p_{\varphi_{n}}$ under $\psi_{n}$. There cannot be polynomial tree-size explicit $\sigma_{n}$-definitions of $p_{\varphi_{n}}$ under $\psi_{n}$ since these would be propositional formulas equivalent to $\varphi_{n}$ of polynomial tree-size. It follows from the proof of Theorem 5.16 that Craig interpolants can also not be of polynomial tree-size.
In what follows, we only consider the dag-size of a formula. So we say that propositional logic has polysize Craig interpolants if there is a polynomial function $f:\mathbb{N}\times\mathbb{N}\rightarrow\mathbb{N}$ such that for any $\varphi,\psi$ with $\varphi\models\psi$ there exists a Craig interpolant $\chi$ for $\varphi,\psi$ with $|\chi|\leq f(|\varphi|,|\psi|)$. We say that propositional logic has polysize uniform interpolants if there is a polynomial function $f:\mathbb{N}\rightarrow\mathbb{N}$ such that for any $\varphi$ and ordering ${\bf p}=p_{1},\ldots,p_{n}$ of any subset of $\mathit{sig}(\varphi)$, there exists a propositional formula $\chi$ equivalent to $\exists{\bf p}.\varphi$ with $|\chi|\leq f(|\varphi|)$, where $\exists{\bf p}.\varphi$ stands for $\exists p_{1}\cdots\exists p_{n}.\varphi$.
@@
Conversely, we show that ${\bf NP}\subseteq{\bf P}_{{\bf/poly}}$ implies that uniform interpolants of polynomial size exist. Let $\exists P_{n}$ denote the set of formulas of the form $\varphi=\exists q_{1}\cdots\exists q_{m}.\psi$ with $\psi$ a propositional formula and $\mathit{sig}(\varphi)=\{p_{1},\ldots,p_{n}\}$ the remaining atoms in $\varphi$ distinct from $q_{1},\ldots,q_{m}$. It suffices to show that ${\bf NP}\subseteq{\bf P}_{{\bf/poly}}$ implies that formulas in $\exists P=\bigcup_{n\in\mathbb{N}}\exists P_{n}$ are equivalent to polysize propositional formulas. Let $w_{\varphi}\in\{0,1\}^{\ast}$ be an encoding of formulas $\varphi\in\exists P$. We regard every $w\in\{0,1\}^{n}$ as a model with $w(p_{i})=w_{i}$ for $1\leq i\leq n$. Consider the set